solve each linear system by subtracting the equations....
1) -x-y=8 2x-y=-1
2) x+y=-7 3x+y=-9
3) x+3y=19 x-y=-1
4) 3x-3y=-15 3x+3y=3
AND MORE TO ASK THANKYOU

Answer by checkley77(12844)

1) -x-y=8
-(2x-y=-1)
----------------
-3x=9
x=9/-3
x=-3 ans.
-(-3)-y=8
3-y=8
-y=8-3
y=-5 ans.
**************************
2) x+y=-7
-(3x+y=-9)
-----------------
-2x=2
x=2/-2
x=-1 ans.
-1+y=-7
y=-7+1
y=-6 ans.
**************************
I'll leave the remaing 2 for you to practice with.
3) x+3y=19 x-y=-1 [ans. x=4 & y=5]
4) 3x-3y=-15 3x+3y=3 [ans. x=-2 & y=3]